OpenStudy (anonymous):

Find the sum of the convergent series 8+6+9/2+27/8+....

OpenStudy (anonymous):

what is the common ratio?

OpenStudy (anonymous):

not given

OpenStudy (anonymous):

i think its geometric but not sure

OpenStudy (anonymous):

Notice that \[8\times\frac{6}{8}=6\] \[6\times\frac{6}{8}=\frac{36}{8}=\frac{9}{2}\] \[\frac{9}{2}\times\frac{6}{8}=\frac{54}{16}=\frac{27}{8}\] What's the common ratio?